### v2

# All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark

Abstract
We introduce MAIA (Multimodal AI Assessment), a native-Italian benchmark designed for fine-grained investigation of the reasoning abilities of visual language models on videos. MAIA differs from other available video benchmarks for its design, its reasoning categories, the metric it uses, and the language and culture of the videos. MAIA evaluates Vision Language Models (VLMs) on two aligned tasks: a visual statement verification task, and an open-ended visual question-answering task, both on the same set of video-related questions. It considers twelve reasoning categories that aim to disentangle language and vision relations by highlighting the role of the visual input. Thanks to its carefully taught design, it evaluates VLMs’ consistency and visually grounded natural language comprehension and generation simultaneously through an aggregated metric revealing low results that highlight models’ fragility. Last but not least, the video collection has been carefully selected to reflect the Italian culture, and the language data are produced by native-speakers. [^1] Data and code will be available upon acceptance.
All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark
Davide Testa1,2, Giovanni Bonetta2, Raffaella Bernardi3, Alessandro Bondielli4,5, Alessandro Lenci5, Alessio Miaschi6, Lucia Passaro4, Bernardo Magnini2 1Università di Roma La Sapienza, 2Fondazione Bruno Kessler (FBK), 3CIMeC, DISI, University of Trento 4Dept. of Computer Science, University of Pisa, 5Dept. of Philology, Literature and Linguistics, University of Pisa 6Istituto di Linguistica Computazionale "A. Zampolli" (CNR-ILC), ItaliaNLP Lab, Pisa davide.testa@uniroma1.it, {dtesta, gbonetta, magnini}@fbk.eu, raffaella.bernardi@unitn.it {name.surname}@unipi.it, alessio.miaschi@ilc.cnr.it
1 Introduction
Figure 1: Structure of the MAIA benchmark. For each video there are two questions related to 12 reasoning categories. For each question there is a pool of 8 answers, which are associated to a true-false statement pair. The example reported in the Figure is related to Out-of-scope reasoning category.
Vision and Language models entered the Computer Vision and NLP scenes more than a decade ago pushed by theoretical (e.g., Baroni, 2015) and application-oriented  (e.g., Bigham et al., 2010) motivations. Their success on combining image and text has been monitored and summarized in various surveys, from the earlier ones on Visual Question Answering (Bernardi and Pezzelle, 2021) to the more recent ones focusing on Visually grounded LLMs (e.g., Caffagni et al., 2024; Li et al., 2024b). Researchers have always felt the need to target Vision and Language Models (VLMs) shortcomings, developing carefully designed benchmarks consisting of a suit of VL tasks to evaluate a variety of capabilities (Kushal et al., 2019). The minimal pair task, contrasting a caption with its foil (Shekhar et al., 2017) has been applied to large-scale linguistic phenomena Parcalabescu et al. (2022) and extended to highlight weakness of VLMs on Video QA Kesen et al. (2023). This trend focuses on visually grounded natural language understanding Kesen et al. (2023).
Today VLMs are trained to generate text and are known to excel at it. We argue that the evaluation of Natural Language Generation (NLG) and Natural Language Understanding (NLU) competence should always be pursued together: An agent that can answer questions about an event must understand it too. Yet, existing benchmarks tend to treat comprehension and generation separately, often relying on independent datasets and evaluation protocols. This fragmented design prevents an assessment of a model’s robustness, in other words, its ability both to understand and generate visually grounded text. Moreover, success in NLU should be claimed only through evaluation regimes that monitor models’ consistency across answers, showing they are insensitive to surface variations.
To address these limitations, we present MAIA: a competence-oriented benchmark consisting of two paired tasks – multiple-choice Visual Statement Verification (VSV) and Open-Ended Visual Question Answering (OEVQA) – containing aligned datapoints. For instance, in the example in Figure 1, for the video showing a pizza cooking in a wood-fired oven, the VSV contains the Truth and False statements (TS and TF) A: There is no cake in the video and B: There is a cake in the video, and the OEVQA contains the (aligned) question What is the cake made of?. Such interleaved data let us evaluate models’ robustness: we evaluate a model positively, only if it performs correctly both on the visual statement verification task (NLU) – it selects There is no cake in the video – and on the Open-ended VQA (NLG) by generating something like There is a pizza, not a cake; while we evaluate it negatively, if it performs correctly only on one of the two tasks. Moreover, the VSV is organized in pools containing 8 TS-FS pairs that differ only on the surface, letting us evaluate the model’s consistency. We categorize questions based on the reasoning they elicit; for instance, the question in Figure 1 is "out-of-scope". MAIA spans twelve reasoning categories, helping highlight the role of language and visual modalities across them. Finally, it implements an all-in-one evaluation philosophy that lets us evaluate both models’ robustness and consistency through its Aggregate Metric. Last but not least, MAIA is based on native Italian videos. To the best of our knowledge, this is the first benchmark for the Italian language on videos.
Contributions.
In the paper, we: (i) present MAIA, the first Italian benchmark specifically designed to assess the reasoning abilities of VLMs on video data; (ii) evaluate multiple VLMs, highlighting how their performance and their reliance on linguistic or visual cues varies across reasoning categories; (iii) demonstrate the importance of evaluating models from multiple perspectives, not only in terms of their competencies, but also their robustness and consistency; (iv) propose a novel metric for evaluating visually grounded comprehension and generation simultaneously.
2 Related Work
Diagnostic benchmarks for VLMs.
Various types of benchmarks have been proposed since the raise of VLMs. From the single task-oriented benchmarks (e.g. Antol et al. (2015); Das et al. (2017)), attention has now moved to task collections Xu et al. (2024); Lee et al. (2024b) in which models show impressive performance. As in the early phase Johnson et al. (2017); Shekhar et al. (2017); Suhr et al. (2017), such success is mitigated by the use of diagnostic benchmarks, such as Parcalabescu et al. (2022); Thrush et al. (2022); Chen et al. (2023); Bugliarello et al. (2023); Bianchi et al. (2024) and carefully curated benchmarks such as Xiao et al. (2024); Tong et al. (2024). The third type of benchmarks available focus on the VLMs competence in a holistic fashion, evaluating advanced perception and reasoning with domain-specific knowledge Yue et al. (2024b). A similar picture emerges for video-based VLMs. Here as well, early surveys call for careful evaluation (e.g. Zhong et al. (2022)), task-oriented benchmarks show impressive performance Grunde-McLaughlin et al. (2021); Yu et al. (2023), while fine-grained ones pinpoint important weaknesses Kesen et al. (2023), and competence-based analysis highlight there is significant room for improvement in multimodal video understanding (Patraucean et al., 2023). Finally, both Tong et al. (2024) for images and Kesen et al. (2023) for videos manage to highlight VLMs shortcomings by imposing a more stringent task-accuracy metric that account for model consistency across very similar data or correlated competencies. Thanks to the richness of MAIA data collection, we adopt such severe, and hence robust, evaluation code and propose a novel aggregate metric. Building on prior work, MAIA targets a low-resource language and the underexplored video domain. While some benchmarks Das et al. (2024); Zhang et al. (2023) include limited Italian multiple-choice tasks, none focus on high-level reasoning in Italian video contexts or analyze distributional biases in VLMs.
Video Reasoning Benchmarks.
Widely used benchmarks, such as AGQA Grunde-McLaughlin et al. (2021) and MVBench Li et al. (2024c) focus on explicit visual elements (e.g., entity, action, and the spatio-temporal reasoning involving them), instead MAIA’s categories focus on the interplay between language and vision, especially when this relation is implicit or must be inferred, a dimension largely neglected in prior Video QA benchmarks. Yue et al. (2024a) includes multiple-choice and open-ended data points from entirely independent data sets with different origins and content, and reports the average performance across distinct tasks. Instead, MAIA’s NLU and NLG data points are aligned, a unique feature of MAIA framework, as such it introduce an Aggregate Accuracy metric specifically designed to ensure that the model’s performance is evaluated consistently across multiple-choice and open-ended questions derived from the same underlying data. There are few other video-text benchmarks including both these formats (e.g., Zhou et al. (2025); Peng et al. (2024)), however, to the best of our knowledge, none of them attempt to define a unifying metric, as we do in MAIA.
3 The MAIA Benchmark
Figure 2: Overview of MAIA reasoning categories (we illustrate 9/12 categories, see the Appendix for the examples about the missing ones). For each of the 100 videos, it contains 2 questions for each of the 12 categories; for each question, it has 8 answers, and each of these answers has a corresponding True and False statement pair.
MAIA (Multimodal AI Assessment) is an evaluation framework designed to assess the reasoning capabilities of VLMs in video-based contexts.
3.1 Dataset
We outline the steps for collecting, generating, and validating the MAIA dataset. Validation steps consists in a qualitative analysis and revision of the data, when necessary. [^2] Examples are in English for readability. Appendix A provides more details, Italian examples, and translations.
Video Collection. We gathered 100 short (ca. $30$s) videos from YouTube Italy. The selection covers various aspects of Italian culture, including cities, art, food, sports, and daily activities (e.g., cooking pasta, having coffee, or watching a soccer match). Preference was given to videos featuring people and close-up shots. An automated script retrieved videos using thematic keywords and ensured Creative Commons compliance.
Reasoning Categories. We defined 12 reasoning categories aiming to probe the cognitive and linguistic skills of multimodal models and to explore the relation between language and vision, while forming the core of the benchmark for evaluating reasoning and grounding in an Italian context.
Questions and Answers Collection. The annotation process was carried out in two phases. In the first phase (question creation), $12$ qualified annotators wrote $2$ open-ended questions [^3] Yes/No and audio-based questions were prohibited. per video for each category, ensuring diversity in entities and events. [^4] Annotators were paid €100 for their work. A manual review verified adherence to guidelines and semantic categories. In the second phase (answer collection), we used Prolific [^5] https://www.prolific.com/ to solve the task, targeting Italian-native participants with specific cultural criteria (Appendix A.3). Each annotator answered 12 out of 24 questions per video [^6] Annotators were paid £7 per hour., focusing on detailed, visually grounded responses. Each question was answered by eight annotators, resulting in $19,200$ responses. Two semi-automatic validation checks were applied to the output: ($1$) semantic consistency with the corresponding question, and ($2$) contradiction tests within each answer pool (Appendix A.3) [^7] We found that 90.25% of the 8-answer pools exhibit full agreement, as they do not contain any contradictions. The remaining 9.75% was manually reviewed by an additional annotator to resolve inconsistencies. This shows agreement among annotators, that cannot be measured through standard annotator agreement metrics due to the open-ended nature of the task.. Through a post-processing of the lexicon, we made sure that the final 8-answer pools are lexical diverse. [^8] The lexical overlap within the 8-answer pool is $21.90$%, similarly to the overlap between pairs randomly extracted from each pool ($22.21$%).
Statement Collection. As shown in Figure 2, TSs are descriptive declarative sentences that accurately align with the visual content of videos. TSs describe videos from different semantic perspectives, according to MAIA semantic categories. TSs were generated using GPT-4o (prompts are reported in Figure 12 of the Appendix): for each question, it is given the 8 human generated answers and it is prompted to produce 8 TSs by combining the content of the question with the one of the corresponding answer. Again, post-processing techniques ensured high lexical variability within each pool reducing lexical overlap. FSs are incorrect descriptions automatically generated by prompting GPT-4o and created by modifying elements related to a reasoning category while maintaining the sentence structure (Appendix A.7). FSs were validated through two semi-automatic checks: a check to ensure minimal modification while maintaining incorrectness for the semantic category, and a test to confirm that FSs contradict their corresponding TSs (Appendix A.8). This process produced $19,200$ high-quality FSs aligned with their corresponding TSs.
3.2 Reasoning Categories
We report the reasoning categories in MAIA. Figure 2 provides examples of aligned question, answer, TS, and FS. More details in Appendix A.2.
Causal.
Focuses on questions about the cause or effect of an event. It provides a comprehensive test of a model’s ability to infer and describe causality within events. It can address either explicit (observable in the video) or implicit (inferred from the visible effect) causes/effects.
Counterfactual.
Focuses on hypothetical events that do not occur in the video but could happen under certain conditions. It tests a model’s ability to reason about plausible scenarios grounded in the video’s context.
Implicit.
Involves questions about entities or events that are either not explicitly visible in the video (Total Implicit) or no longer visible (Partial Implicit), but can still be reasonably inferred. It evaluates a model’s ability to deduce implicit details based on context.
Out-of-scope.
Assumes the presence of entities or events not actually shown in the video, asking about properties of these nonexistent elements. It tests the model’s ability to handle multimodal hallucinations and its tendency to make assertive, yet incorrect, responses.
Planning.
Inquires about the sequence of actions needed to achieve a specific goal related to the video. It assesses the model’s ability to infer and plan the necessary steps based on the visual cues.
Sentiment.
Focuses on sentiment, mood, attitude, or emotions of characters towards other entities or events in the video. It evaluates the model’s ability to recognize and identify the emotional cues.
Spatial.
Focuses on the location of entities in space, either applicable to the entire video (Total Spatial) or specific moments and events (Partial Spatial). It assesses the model’s ability to infer stable and time-dependent spatial relationships, determine relative positioning, and demonstrate grounding competencies.
Temporal.
Relies on when something happens, either in relation to other events (Partial Temporal) or the duration of an event (Duration). It evaluates the model’s ability to infer temporal relationships, event sequences, and durations from visual content.
Uncertainty.
Arises when insufficient information is provided in the video to give a precise answer. It tests model’s ability to handle situations with ambiguous or incomplete information, assessing its tendency to make assertive (rather than uncertain) responses.
4 Experimental Setting
We run several experiments to test modern VLMs on the MAIA benchmark in a zero-shot setting. To capture different VLM behaviors, strengths and limitations, we defined two tasks, aligned on the same datapoints: a multiple choice task, Visual Statement Verification (VSV), and a generative task, Open-ended Visual Question Answering (OEVQA).
4.1 Task1: Visual Statement Verification
VSV is a multiple-choice task where a model is presented with a true-false statement pair related to a MAIA question (see section 3.1) for a given video, and has to select the true statement. The two statements are randomly assigned to two labels, A and B, and the model is asked to generate only the label. We chose the prompt through an extensive evaluation of $32$ variants ($16$ in Italian and $16$ in English), with the best-performing Italian prompt ultimately selected. Performance for VSV is measured with accuracy, i.e., the proportion of correctly selected true statements over the total statement pairs.
4.2 Task2: Open-ended VQA
OEVQA is a generative task, where models are tested on their ability to provide correct open-ended answers to a question related to video content. The model receives as input a prompt question and a video, and is tasked with generating a correct answer. The prompt used in the experiments was selected as the best-performing among $10$ tested variants ($5$ in Italian and $5$ in English). Generated responses are then evaluated according to two approaches: similarity-based and LLM as a judge.
Similarity-based metrics
compare a response against the pool of 8 reference answers available in MAIA. We used five token-level metrics: ROUGE Lin (2004), BLEU Papineni et al. (2002), BERT-Score Zhang et al. (2020), METEOR Lavie and Agarwal (2007) and CIDEr Vedantam et al. (2015).
LLM as a judge.
While similarity metrics help rank VLMs, they fail in assessing answer correctness. To address this, we adopt an LLM-as-a-judge approach Gu et al. (2025), using GPT-4o to evaluate whether an answer is semantically consistent with at least one of the eight MAIA references, prioritizing meaning over surface-level structure (Appendix B.3). Following Bavaresco et al. (2024), we validate this method on 100 samples: annotations by two human raters and GPT-4o show strong agreement, with a Fleiss’ Kappa of 0.82.
(a) Task 1: VSV
(b) Task 1: VSV (pool)
(c) Task 2: OEVQA
Figure 3: Qwen2.5-VL’s fingerprint through MAIA’s reasoning categories: (a) illustrates model performance in NLU, Task 1, when TS-FS pairs are independent, while (b) reports performance on the same task with TS-FS pairs aggregated within their 8-items pool, penalizing models’ inconsistency; (c) visualizes the performance on NLG, Task 2.
Models Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Partial Total Partial Total Duration Partial
Unimodal 0.05 0.01 0.12 0.04 0.04 0.02 0.17 0.04 0.04 0.01 0.01 0.03 0.04
Black video InternVL2 0.18 0.07 0.40 0.43 0.08 0.06 0.73 0.03 0.03 0.03 0.02 0.20 0.05
Llava-Next-Video 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.01 0.01 0.00 0.00 0.01 0.01
Llava-oneVision 0.14 0.01 0.04 0.59 0.04 0.00 0.83 0.04 0.04 0.01 0.01 0.06 0.01
Qwen-2.5-VL 0.22 0.11 0.36 0.41 0.17 0.16 0.87 0.04 0.06 0.09 0.07 0.15 0.16
1-Frame InternVL2 0.28 0.32 0.41 0.32 0.11 0.33 0.47 0.17 0.33 0.15 0.39 0.11 0.21
Llava-Next-Video 0.08 0.12 0.23 0.01 0.03 0.21 0.01 0.06 0.11 0.03 0.14 0.01 0.03
Llava-oneVision 0.32 0.35 0.29 0.64 0.14 0.32 0.65 0.21 0.35 0.12 0.36 0.15 0.20
Qwen-2.5-VL 0.36 0.36 0.28 0.64 0.16 0.44 0.81 0.25 0.39 0.14 0.39 0.21 0.24
32-Frames InternVL2 0.31 0.41 0.35 0.38 0.12 0.42 0.39 0.28 0.38 0.18 0.43 0.11 0.30
Llava-Next-Video 0.03 0.04 0.04 0.01 0.03 0.09 0.00 0.01 0.06 0.01 0.04 0.01 0.01
Llava-oneVision 0.38 0.51 0.21 0.61 0.19 0.51 0.45 0.39 0.47 0.26 0.48 0.11 0.33
Qwen-2.5-VL 0.44 0.53 0.29 0.63 0.23 0.50 0.67 0.43 0.56 0.28 0.55 0.28 0.32
Table 1: VSV (Task $1$): aggregated accuracy of correct pools (8/8) across reasoning categories, penalizing models’ inconsistency.
4.3 Baselines
We implemented three baselines for our tasks.
Unimodal.
This baseline applies only to Task 1 and selects the most probable statement in a TS-FS pair. It serves as a unimodal language baseline, reflecting the distributional biases of LLMs. Probabilities of TS and FS are first estimated on five open-weight LLMs that have shown good performance on a variety of Italian tasks Magnini et al. (2025): Llama-3.1 ($8$B-Instruct), LLaMAntino-2 ($7$B), LLaMAntino-3-ANITA ($8$B-Instruct), Gemma ($7$B) and Qwen2.5 ($7$B-Instruct). For each TS-FS pair, we selected the item with the highest probability among the five models.
Black video.
replaces MAIA videos with a fully black clip, used as a proxy for a no-video condition This setup minimizes access to visual features, pushing the model to rely mainly on the language component, while the true unimodal evaluation remains the previous one.
1-Frame.
This baseline considers only the first frame for each MAIA video, this way reducing the capacity of a VLM to capture visual features and facing the one-frame “static appearance bias” (Lei et al., 2023).
4.4 Vision-Language-Models.
We benchmarked four recent VLMs sourced from the Hugging Face Hub, representing state-of-the-art approaches in Vision-Language tasks: [^9] More details about VLMs are reported in Appendix B.1 InternVL2 (8B, Chen et al. (2024)), LLaVA-NeXT-Video (7B, Zhang et al. (2024b)), LLaVa-OneVision (7B, Li et al. (2024a)), and Qwen2.5-VL (7B, Qwen et al. (2025)). All models accept a [video, text] pair as input, and uniformly sample 32 frames from the video.
5 Results
This section reports the results obtained in our experiments for Task 1 and Task 2 independently.
5.1 Results on Visual Statement Verification
Table 1 shows VLM accuracy across the three settings (black video, 1-Frame, 32-Frames) and reasoning categories when models consistently answers correctly (i.e., by choosing 8/8 the TSs within the pool). Qwen2.5-VL achieves the highest accuracy in the 32-Frame setup with an average score of 0.44, marking an 8-point improvement over the correspondent 1-Frame settting. Llava-Next-Video shows the weakest performance across all three configurations, likely due to its underlying Vicuna-7B LLM, weakly trained in Italian. All models outperform the unimodal baseline (0.05): in the 32-Frame setting, gains go from +31 points (InternVL2) to +44 (Qwen2.5-VL), confirming the use of visual cues to counteract language-driven biases.
Figure 3 helps visualizing the comparison across reasoning categories. Here we report results only for Qwen-2.5-VL, the best model, while similar figures for the other models are in Appendix B (Figure 15, 16 and 19). The most difficult categories are Counterfactual, Planning, Implicit Partial, Spatial Partial, and Temporal-Duration. On the other side, Causal, Sentiment, Implicit Partial and Total, and Spatial Partial and Total are the categories for which the model profits the most from visual clues, by leveraging the broader visual window provided in the 32-Frame setting.
Models’ consistency.
Figure 3a vs.  3b highlights the role of the consistency check in MAIA by using a pool of 8 TS-FS pairs. When pairs are considered independently, as it is usually done in VLM benchmarks, the model’s performance increases significantly, showing that it relies on spurious correlation, effects that is strongly mitigated by the MAIA’s severe evaluation regime. Such trend is even much more visible in the black video setting than in the 32-Frames one (e.g. +51 vs. +40 for Qwen2.5-VL). By systematically comparing models’ performance in the independent and pool-aggregated settings, we see that this is a general finding across models. Moreover, we find that Qwen-2.5-VL is not only the best-performing model, but also the most consistent, with a lower drop in accuracy, particularly in the 32-Frames setting (Details in Appendix B.2).
5.2 Results on Open-ended Generation
Models Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Partial Total Partial Total Duration Partial
Black video InternVL2 0.37 0.42 0.60 0.30 0.68 0.43 0.08 0.21 0.23 0.36 0.25 0.55 0.25
Llava-Next-Video 0.27 0.43 0.47 0.30 0.40 0.33 0.51 0.21 0.12 0.16 0.04 0.12 0.16
Llava-oneVision 0.40 0.66 0.68 0.29 0.60 0.60 0.28 0.26 0.24 0.36 0.18 0.37 0.23
Qwen-2.5-VL 0.35 0.36 0.69 0.08 0.54 0.74 0.23 0.24 0.19 0.30 0.19 0.41 0.20
1-Frame InternVL2 0.44 0.57 0.65 0.21 0.65 0.60 0.10 0.33 0.38 0.39 0.47 0.53 0.35
Llava-Next-Video 0.32 0.30 0.56 0.20 0.46 0.60 0.29 0.18 0.32 0.24 0.27 0.20 0.19
Llava-oneVision 0.50 0.59 0.78 0.15 0.66 0.74 0.37 0.39 0.54 0.39 0.51 0.54 0.33
Qwen-2.5-VL 0.51 0.56 0.76 0.32 0.60 0.79 0.40 0.35 0.50 0.38 0.53 0.55 0.38
32-Frames InternVL2 0.49 0.54 0.68 0.28 0.64 0.62 0.11 0.45 0.48 0.47 0.51 0.57 0.46
Llava-Next-Video 0.33 0.37 0.38 0.16 0.42 0.48 0.27 0.24 0.37 0.29 0.39 0.32 0.29
Llava-oneVision 0.53 0.67 0.79 0.11 0.65 0.79 0.23 0.55 0.56 0.51 0.62 0.40 0.46
Qwen-2.5-VL 0.61 0.71 0.80 0.43 0.60 0.85 0.55 0.55 0.60 0.50 0.70 0.54 0.53
Table 2: OEVQA (Task $2$): accuracy of correct answers with LLM-as-a-judge.
Model Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Partial Total Partial Total Duration Partial
Black video Qwen-2.5-VL 0.09 0.05 0.29 0.05 0.09 0.14 0.02 0.02 0.20 0.01 0.03 0.10 0.04
1-Frame Qwen-2.5-VL 0.22 0.25 0.23 0.25 0.11 0.40 0.33 0.16 0.28 0.06 0.28 0.17 0.14
32-Frames Qwen-2.5-VL 0.31 0.41 0.26 0.31 0.16 0.48 0.41 0.31 0.38 0.19 0.44 0.22 0.19
Table 3: Aggregate accuracy on Task $1$ (NLU) and Task $2$ (NLG) (consistency and robustness) on Qwen-2.5-VL across reasoning categories.
Table 2 reports the results on Task 2. The best performing model is again Qwen2.5-VL, reaching 0.61 accuracy in the 32-Frames setting. Hardest categories are Sentiment, Implicit Partial, Spatial Partial, Temporal, and Planning. Figure 3c shows that model profits from the 32 frames for all categories, and poorly performs when provided only with black videos, in almost all cases. Surprisingly, for Sentiment and Counterfactual, a relatively high accuracy is obtained already with black videos. Interestingly, differences also emerge when comparing the 1-Frame vs. 32-Frames settings. For Temporal Partial, and Implicit Partial, and for Spatial Total, the 1-Frame is not enough, while 32 frames increase performance. This difference is less pronounced or does not hold for Temporal Duration, Counterfactual, Planning, and Sentiment. For example, in the Implicit Partial category, the 32-Frame model answers correctly to How the man in the flannel shirt wears his hair? with The man in the flannel shirt has shaved head, while the 1-Frame setting hallucinates with The man in the flannel shirt keeps his hair in a braid.
Models ROUGE BertScore BLEU METEOR CIDEr
InternVL2 0.61 0.84 0.38 0.59 1.18
LLaVa-NeXT-Video 0.46 0.79 0.21 0.45 0.65
LLava-oneVision 0.58 0.83 0.40 0.55 1.08
Qwen-2.5-VL 0.58 0.83 0.38 0.61 0.98
Table 4: VLM performance for OEVQA (Task 2) according to similarity-based metrics.
Table 4 highlights how similarity-based metrics (e.g., ROUGE, BLEU) often do not align with semantic correctness. InternVL2 scores highest on surface-level similarity (e.g., ROUGE:$0.61$, BERTScore:$0.84$) but lower when considering the LLM-as-a-judge metric ($0.49$), while Qwen2.5-VL offers more (semantic) accurate answers. LLaVA-NeXT-Video underperforms across all evaluations.
5.3 Discussion
The results presented for Task 1 and 2 clearly highlight two key findings. First, within each task, the role of the information extracted from videos is unequally distributed across reasoning categories. The star-shaped Figure 3b illustrates such differences: the star’s picks highlight the categories that profit from the visual input the most: Spatial Total and Partial, Total Implicit, Causal, Out-of-scope, and Sentiment. On the other hand, and quite surprisingly, Qwen2.5-VL handles Counterfactual and Uncertainty better when provided the black videos, which are expected to be uninformative, than with the full 32 frames. This calls for a deeper analysis of the reason behind such a result. Overall, this shows that MAIA categories are extremely useful to factorize the contribution of the visual vs. linguistic components of VLMs, favoring a more nuanced analysis of their actual abilities.
A second noteworthy fact is the effect of the task design. By comparing Table 1 and Table 2, we see that overall accuracy is higher in Task 2 than in Task 1, even though the underlying information the model has to exploit to perform the tasks is the same, given the alignment between their data points. Such an increase is found across models: +0.2 increase on average – in line with the Generative AI Paradox – they are better at generating than at understanding text West et al. (2024). Figure 3b vs. Figure 3b illustrates such a difference for Qwen2.5-VL. Interestingly, the difference is more pronounced for some categories, as shown by the fact that Figure 3c no longer has the shape of a star; for instance, Planning, Counterfactual, and Spatial Partial improve the most. On the other hand, Uncertainty and Out-of-scope accuracy drops in the NLG task; from a qualitative analysis, we saw that this is mostly due to the generation of hallucinations (Appendix B.4).
6 Aggregating Understanding and Generation
A distinctive feature of MAIA is the alignment between the VSV and OEVQA tasks: both are grounded in the same question and the same video. While Section 5 reported their results separately, we now combine them into a unified evaluation framework that jointly tests comprehension and generation abilities in VLMs. We introduce Aggregate Accuracy (Agg-Acc), a metric rewarding models that: (i) consistently select the correct statement (TS) over all 8 TS-FS pairs in Task 1, and (ii) generate a correct answer to the same question in Task 2, according to our LLM-as-a-judge evaluation. The idea is to capture overlapping abilities (i.e. knowledge and reasoning) for the two aligned tasks, evaluating models’ robustness. Aggregate Accuracy is defined as follow:
$Agg\text{-}Acc(M,q)=\begin{cases}1&\text{if }\forall(TS,FS)\in S_{q},\\ &a_{M}(TS,FS)=TS\\ &\text{and }a_{M}(q)\text{ is correct}\\ 0&\text{otherwise}\end{cases}$
where $M$ is the model, $q$ a question, $S_{q}$ the set of TS-FS pairs, and $a_{M}$ the model’s answers. Intuitively, given a question $q$ on a video $v$, we reward the ability of a VLM to both select a correct answer $TF$ from a $TF-FS$ pair related to $q$, and to generate a correct answer $a$ to $q$, as discussed in Section 1 when commenting the question What is the cake made of? and the aligned data of the NLU and NLG tasks in Figure 1.
Table 3 reports Agg-Acc for our best model Qwen2.5-VL: 0.31 with 32-frames, 0.22 with 1-frame, and only 0.09 with black video; this shows the more challenging nature of the aggregate task, and that a single frame is not that robust, except for Counterfactual where reasonable answers can be produced without visual features. Planning, Spatial Partial, and Temporal Duration remain challenging even with 32-Frames. Results confirm that MAIA’s aggregated understanding&generation task creates a harder benchmark, laying the ground toward a more objective VLM evaluation.
7 Conclusion
We introduce MAIA, a benchmark designed for fine-grained investigation of the reasoning abilities of VLMs on videos. MAIA has two aligned tasks: a visual statement verification task (NLU), and a open-ended visual question answering task (NLG), both on the same set of video related questions. First, we provided a in-depth analysis of the two tasks independently, showing the importance of evaluating model with answer-pools to account for model consistency. Then, we couple comprehension and generation in an aggregated evaluation framework, arguing that the aggregated "all-in-one" understanding&generation task is a challenging and natural setting toward a more objective VLM evaluation. As for the future, it would be interesting to see whether our framework promote models that undergo learning paradigms tightly integrating these two capabilities, as in Gul and Artzi (2024).
Limitations and Future Directions
We acknowledge that the number of videos in MAIA may seem relatively small, with only 100 samples. However, their combination with 12 reasoning categories results in 19,200 samples for Task 1 and an equal number of question-answer pairs, providing a robust evaluation set, not intended for any kind of training. Still, these 100 videos constitute only the initial core of a broader evaluation framework planned for future development. Moreover, we are aware that our most probable baseline, constructed using probabilities derived from our set of LLM’s logits, poses a limitation that we intend to address in future work. We plan to compare each VLM with its corresponding LLM to obtain more reliable results for proper comparisons and analyses of potential statistical biases. Regarding Task $2$, we aim to dive deep into the comparison between the similarity metrics used and the emerging topic of LLMs as judges. Additionally, we intend to further investigate this latter approach to assess its actual validity as a reliable evaluation method. This direction will help us refine accuracy metrics, ultimately enhancing our ability to rigorously test model robustness and consistency across our two specular tasks. Furthermore, we acknowledge that our evaluation did not include large-scale propri- etary models (e.g., ChatGPT). Our focus was pri- marily on testing the performance of open-source and easily exploitable language models to provide a comprehensive overview of their capabilities on the benchmark. Nevertheless, extending our evalu- ation to bigger models would be beneficial to gain a broader understanding of their performance and to contextualize our findings within the wider land- scape of LLM research. Finally, we are also aware that our benchmark has not yet been compared with existing ones to assess its relative difficulty and challenge level. This limitation comes from the lack of comparable resources in Italian. As a future direction, we plan to translate MAIA into English and replicate our experiments, enabling more meaningful comparisons with similar English-language benchmarks.
Acknowledgments
This work has been carried out while Davide Testa was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Fondazione Bruno Kessler (FBK). Giovanni Bonetta and Bernardo Magnini were supported by the PNRR MUR project PE0000013-FAIR (Spoke 2). Alessandro Lenci was supported by the PNRR MUR project PE0000013-FAIR (Spoke 1). Alessio Miaschi was supported by the PNRR MUR project PE0000013-FAIR (Spoke 5). Lucia Passaro was supported by the EU EIC project EMERGE (Grant No. 101070918). Alessandro Bondielli was supported by the Italian Ministry of University and Research (MUR) in the framework of the PON 2014-2021 “Research and Innovation" resources – Innovation Action - DM MUR 1062/2021 - Title of the Research: “Modelli semantici multimodali per l’industria 4.0 e le digital humanities.”
Appendix A MAIA: Benchmark Creation Process
This appendix provides a detailed explanation of the step-by-step process involved in developing the dataset of the benchmark. Each sub-section corresponds to a specific stage, outlining the methodologies, decisions, and criteria that guided the creation of MAIA as an evaluation framework. Figure 4 illustrates the underlying logic of the evaluation framework, outlining its construction process and the tasks involved. Similarly, Table 5 shows MAIA’s dataset structure after the creation and validation process which is accurately described in the following Section. The final structure of the dataset is shown in Figure 5, which accurately provides examples of aligned question, answer, TS, and FS.
Figure 4: Workflow of the MAIA evaluation framework, integrating dataset construction with its application to the two main benchmark tasks.
Feature n
Videos $100$
Semantic Categories 1212121211footnotemark: 1 ($9$ Macro-Cat.)
Questions $2,400$ ($2$ x Category)
8-Answers Pool $2,400$
Answers $19,200$
True Statements $19,200$
False Statements $19,200$
Table 5: MAIA Statistics
Figure 5: Overview of MAIA reasoning categories. For each of the 100 videos, it contains 2 questions for each of the 12 categories; for each question, it has 8 answers, and each of these answers has a corresponding True and False statement pair.
A.1 Video Selection
The first step in creating the benchmark involved the selection of a culturally representative dataset of $100$ short videos coming from YouTube, each lasting approximately $30$ seconds. To reflect Italian culture authentically and to ensure richness and variety, the selection process tried to be focused on the following thematic areas:
  • Locations
  • Food
  • Sport
  • Job
  • Nature
  • Activities
Such topics allowed us to collect a dataset showing locations (indoor and outdoor), iconic Italian cities, and daily activities such as crossing the street, enjoying breakfast at a café, cooking pasta, attending a soccer match or even important events such as Italian local festivals or typical Italian weddings. Priority was given to videos featuring people, particularly with close-up shots, to enhance visual and contextual relevance.
The video collection process was facilitated by an automated script that scraped videos from YouTube matching specific keywords tied to these thematic areas. Additional filters ensured that the content came from YouTube Italy and respected the Creative Commons licensing for the content use and reproducibility. Once the videos were identified, as many of them were too long, the most engaging $30$-second segments were extracted and cut to create a uniform set of $100$ short clips, representing the base of our Italian-native dataset.
A.2 Semantic Categories Definition
This step focused on establishing and refining the semantic categories and phenomena that would form the foundation for analyzing the previously selected videos. This process aimed to ensure that the benchmark comprehensively addressed a wide range of linguistic reasoning categories, providing a robust framework for evaluating vision-language models. Each of these categories was carefully tailored to align with the cultural and contextual richness of the selected video dataset and resulted from a long period of discussion and refinements. Initial categories were proposed and systematically tested against sample videos to assess their applicability and coverage. Feedback from these tests was used to refine definitions, ensuring clarity and reducing overlap between categories. The final set of semantic categories represents a broad spectrum of linguistic and semantic reasoning phenomena, designed to challenge VLMs in interpreting the intricate interaction between visual and linguistic cues in diverse, real-world contexts. These categories form the backbone of the benchmark, allowing the models’ reasoning capabilities to be extensively tested on culturally representative Italian scenes. Next, the following paragraphs present in detail the nine final macro-categories, along with their definitions and some examples of usage. These categories and descriptions were also used to administer the Google forms for collecting data from the annotators.
Causal
This category [^10] Note that in MAIA there are four macro-categories with two fine-grained specifications (i.e., subcategories). The only exception is the Causal category, in which explicit and implicit items are equally represented (100 each). However, we do not consider them subcategories in the same way as the others, since in those cases, the subcategories express entirely different aspects of the same domain. includes two subtypes: Implicit Causal and Explicit Causal, both aimed at reasoning about the causes or effects of events depicted in the video. Thus, it includes reasoning tasks involving both visible and inferred causal relationships, offering a comprehensive test of a model’s ability to infer and describe causality within events.
  • Implicit Causal: This type of question targets the inferred cause of an event, object, or human action visible in the video. The focus is on an implicit cause that cannot be directly observed but must be deduced from the effect presented in the scene. Typical responses involve a logical inference explaining the implicit cause behind the visible effect.
    Example: Suppose a video shows a person at home grabbing an umbrella while going out.
    Italian:
    Q: Per quale motivo la persona prende l’ombrello?
    A: Perchè potrebbe piovere fuori.
    English:
    Q: Why does the person take the umbrella?
    A: Because it might be raining outside.
    In this example, the action of grabbing the umbrella is visible, but the reason (bad weather) is not explicit in the video and must be inferred.
  • Explicit Causal: This type of question addresses direct cause-and-effect relationships visible within the video. The focus is on identifying a specific event, object, or human action (the cause) that led to another event, situation, or state (the effect) or vice versa. Typical responses clearly describe either the cause or the effect based on what is directly observable in the video.
    Example: Suppose a video shows an angry person throwing a glass on the floor, which subsequently shatters.
    Italian:
    Q: Perchè il bicchiere si è rotto?
    A: Perchè la persona lo ha gettato a terra.
    English:
    Q: Why did the glass break?
    A: Because the person threw it on the ground.
    Here, both the cause (throwing the glass) and the effect (the glass breaking) are visible in the video and can be used to provide a direct response.
Counterfactual
This category focuses on questions about hypothetical scenarios that do not actually occur in the video but could take place under specific conditions. These questions explore the consequences of an event or situation that might happen in the video if a certain condition were met. A key requirement is that the hypothetical condition must be based on entities or events visible in the video. Consequently, this category tests a model’s ability to reason about hypothetical scenarios grounded in the context of the video while deriving logical and plausible outcomes from such scenarios.
Example: Suppose a video shows an outdoor concert.
Italian:
Q: Cosa succederebbe al concerto se arrivasse un forte temporale?
A: Il concerto verrebbe interrotto all’istante.
English:
Q: What would happen to the concert if a violent thunderstorm started?
A: The concert would be immediately interrupted.
In this example, the focus of the question (the concert) is visible in the video, while the condition (a violent thunderstorm) is not. The consequence (the concert being interrupted) is not shown in the video but can be reasonably inferred.
Implicit
The implicit category includes questions about entities, events, or their attributes that are not explicitly visible in the video. However, their presence or properties can be reasonably inferred from the context. This category evaluates a model’s ability to infer implicit details based on context, whether the target information was never shown or was previously visible but later obscured.
  • Total Implicit: These questions focus on entities or events that are never directly visible in the video but can be inferred from observable details. A typical answer provides the requested information based on logical inference.
    Example: Suppose a video shows the interior of a house, and suddenly the front door opens, revealing a person soaking wet with a dripping closed umbrella.
    Italian:
    Q: Che tempo fa fuori?
    A: Piove molto forte.
    English:
    Q: What’s the weather like outside?
    A: It’s raining heavily.
    In this case, the focus of the question (the weather outside) is not visible at any point in the video. However, details such as the wet person and dripping umbrella allow for a reasonably confident inference (heavy rain).
  • Partial Implicit: These questions address entities or events that were visible earlier in the video but are no longer visible due to a shift in the scene or because they have moved out of the frame.
    Example: Suppose a video shows a man placing a pen in a drawer and then closing it.
    Italian:
    Q: Dove si trova la penna?
    A: La penna è nel cassetto.
    English:
    Q: Where is the pen?
    A: The pen is inside the drawer.
    In this example, the focus of the question (the pen) is no longer visible in the video. However, earlier information (the man placing the pen in the drawer) allows for a logical and confident answer (the pen is in the drawer).
Out-of-scope
Such a category involves questions about entities or events that are not present in the video at all, asking for properties or details about these non-existent entities or events. A typical response to an out-of-scope question is a negation, stating that the entity or event in question is not present. This category tests the ability of a model to identify and handle irrelevant or non-existent entities within the video content, appropriately responding with a negation when the requested object or event is absent. Thus, it represents an indirect way to test the models on possible multimodal hallucinations and their tendency to be assertive in their responses.
Example: Suppose a video shows a dog and its owner playing in the park, but there are no cars in the scene.
Italian:
Q: Di che colore è la macchina?
A: Non ci sono auto (nella scena).
English:
Q: What color is the car?
A: There is no car (in the scene).
In this example, the focus of the question (the car) is not physically present in the video, nor can its presence be reasonably inferred. When trying to answer the question, no useful information about a car can be found, and the expected response would be a negation, such as "There is no car."
Planning
This category involves questions that request the actions needed to achieve a specific goal related to the video. The typical response to a planning question is a sequence of actions that someone should perform, based on the situation presented in the video, in order to reach the desired outcome. Such a category assesses the model’s ability to infer and plan the necessary steps to accomplish a goal based on the visual cues provided in the video.
Example: Suppose a video shows a dog and its owner playing with a ball in a park, and the owner throws the ball onto a bench.
Italian:
Q: Cosa dovrebbe fare il cane per continuare a giocare col padrone?
A: Dovrebbe correre verso la palla, saltare sulla pacchina, prendere la palla e riportarla al padrone.
English:
Q: What should the dog do to continue playing with its owner?
A: The dog should run toward the ball, jump onto the bench, grab the ball, and bring it back to the owner.
In this example, the focus of the question (the dog) is visible in the video. To answer the question, one can use the information in the video (the ball on the bench) to deduce the series of actions the dog should take (running, jumping, grabbing, and returning the ball) in order to continue the game.
Sentiment
The category involves questions that focus on the sentiment, mood, attitude, or emotion displayed by one or more characters in the video (i.e., animated beings) toward other entities or events in the scene, throughout the entire video. A typical response to a sentiment question may describe a specific sentiment, attitude, or emotion, or it may reflect a neutral stance. This category represents a tool for evaluating model’s ability to recognize and identify the emotional state or attitude of characters based on visual cues, reflecting their reaction or feelings toward the events and other entities in the video.
Example: Suppose a video shows children who appear bored at a birthday party.
Italian:
Q: Che atteggiamento hanno i bimbi?
A: Sono annoiati.
English:
Q: What is the attitude of the children?
A: They are bored.
In this example, the focus of the question (the children) is visible in the video. To answer the question, one can use the visual cues present in the video (expressions and behaviors of the children) to infer the sentiment (boredom) displayed by the characters.
Spatial
Such a category involves questions related to the spatial relationships between entities, objects, or events depicted in the video. It aims at assessing the model’s ability to infer both stable and time-dependent spatial relationships, as well as the ability to determine relative positioning in space and to rely on grounding competencies.
  • Total Spatial: This question asks about the position of entities in space (including their relation to other entities) that remains constant throughout the entire video, disregarding any temporal variations or minimal movements of the entity at different moments in the video. A typical response to this type of question provides specific spatial information valid for the entire duration of the video.
    Example: Suppose a video shows a school lesson with a teacher and students in a classroom.
    Italian:
    Q: Dov’è l’insegnante?
    A: L’insegnante è in classe.
    English:
    Q: Where is the teacher?
    A: The teacher is in the classroom.
    In this example, the focus of the question (the teacher) is visible throughout the video. To answer the question, one can use visible information from the video (classroom, desk) to provide the entity’s spatial position (behind the desk) throughout the video’s duration.
  • Partial Spatial: This question asks about the position of entities in space, but in relation to the time and/or other events occurring in the scene. It may also request the position of one entity relative to another, with a temporal aspect taken into account. A typical response provides spatial information that is specific only to the requested time range in the video.
    Example: Suppose a video shows a school lesson with a teacher and students in a classroom.
    Italian:
    Q: Dove si trova l’insegnante all’inizio del video?
    A: All’inizio del video, l’insegnante è di fronte la cattedra.
    English:
    Q: Where is the teacher at the beginning of the video?
    A: At the beginning of the video, the teacher is standing in front of the desk.
    In this example, the focus of the question (the teacher) is visible in the video. To answer the question, one would use the visual information visible in the specific part of the video (classroom, desk) to provide the spatial position (in front of the desk) relative to the time frame requested (at the beginning of the video).
Temporal
The category includes questions that focus on temporal information. This category studies the model’s ability to infer temporal relationships, sequence of events, and durations from visual content in a coherent manner.
  • Partial Temporal: This question focuses on the temporal properties and relationships of events in the video. The questions may request any type of temporal information about the events or their temporal relationships, except for their duration. For example, asking when something happens or if something happens before or after another event. A typical response provides the event with the specific temporal information requested by the question.
    Example: Suppose a video shows a rock band concert.
    Italian:
    Q: Che succede dopo che il chitarrista inizia a suonare?
    A: Il cantante inizia a cantare.
    English:
    Q: What happens after the guitarist starts playing?
    A: The singer starts singing.
    In this example, the focus of the question (what happens after the guitarist starts playing) is visible at a specific moment in the video. To answer the question, one can use the visible information in that portion of the video (the singer starts singing) to provide the event (the singer starting to sing) as a response.
  • Duration Temporal: This question focuses on a specific property of events in the video: their duration. A typical response provides the specific temporal information required by the question regarding the event’s duration.
    Example: Suppose a video shows a room with a light on and a person switching it off.
    Italian:
    Q: Per quanto tempo la luce rimane accesa?
    A: Per circa 15 secondi.
    English:
    Q: How long was the light on?
    A: For about 15 seconds.
    In this example, the focus of the question (the light on) is visible in the video. To answer the question, one can use the temporal information visible in the video (the person switching the light off) to provide a duration (about 15 seconds) as response.
Uncertainty
This question refers to entities and events that are part of the situation represented in the video, but the scene does not provide enough information to give a precise answer. Therefore, uncertainty questions involve a certain degree of ambiguity in the response, which cannot be fully derived from the video content. The answer may refer to a range of values, state that a precise answer cannot be given, or mention that the answer is a guess and might not be correct. This category tests the model’s ability to recognise and deal with situations in which the available information is insufficient or ambiguous, leading to a response that reflects the uncertainty of the scene and indirectly testing the hypothetical assertive behaviour of such models in answering.
Example: Suppose a video shows a dog.
Italian:
Q: Quanti anni ha il cane?
A: Difficile da dire. / Il cane è probabilmente giovane, ma non si può esserne certi.
English:
Q: How old is the dog?
A: It’s hard to say. / The dog is probably young, but it’s not certain.
In this example, the focus of the question (the dog) is visible in the video. However, if one tries to answer the question, only partial information about the dog’s age is available in the video. As a result, an uncertain answer (e.g., “It’s difficult to tell”) is expected.
A.3 Data collection
This section outlines the data collection process and the methodologies adopted for this step of the research which plays a crucial role in the development of this evaluation framework. This section is divided into two main parts: Questions Collection, which describe the procedure for creating relevant questions based on video content, and Answers Collection, which tells how annotators provided diverse responses to those questions.
Questions Collection
After defining the Semantic Categories on which we based the subsequent analysis, we created $12$ different sets of guidelines (one for each subcategory [^11] The original subcategories were $13$, but due to the difficulty of the causal category, we decided to merge the category of explicit and implicit causals into one group. The category was then assigned to a student who annotated the first $50$ videos with explicit causals and the last $50$ with implicit causals to allow the subcategories to be balanced within the same group.) that were administered to the same number of annotators via Google Forms.
Figure 6: Example of a Google Form used for collecting 2 Questions for each video for each of the assigned category
All the selected annotators were PhD students under $30$ with specializations in Linguistics and Computational Linguistics. They were paid with $100$ euros for the generation task, which was executed in multiple stages through the administration of $1,200$ different forms ($10$ forms per annotator). After carefully reading the rules and instructions in the form, their task was to watch the video as much as they wanted, and to compulsorily generate $2$ questions (i.e., Question_A and Question_B) for each of the $100$ videos in the dataset, as shown in Figure 6 . To ensure variability between the pair of questions about that video, in the second question, the annotators were asked to try as far as possible to change the entities and/or events involved in the first one previously formulated. The final result corresponds to a total of $2,400$ questions ($24$ per video). Figure 7 shows an example of a video with $12$ out of $24$ questions for each semantic category.
Figure 7: Example of a video with its $12$ Questions, each one for a sub-category.
Each form provided contained both the definition of the assigned semantic category with examples (See section A.2), and also general rules to be followed. Each question had to be generated naturally and as an open-ended question. Questions involving a ‘Yes/No’ answer (e.g. ‘Is there a car in the video?’) were not allowed. Finally, for the correct execution of the task, the audio of the video had to be ignored, as the Vision Language Models to be tested could only work on the visual part.
Subsequently, these questions were manually reviewed to ensure quality and category alignment. Thus, while verifying whether the questions generated could actually belong to the specific category and whether they complied with the rules presented in the forms, one third of the generated questions were modified or refined.
Answers Collection
Once the questions were obtained and checked, they were used to set up the answer collection process through Prolific, a platform designed to help researchers recruit participants for high-quality data collection. To follow the initial goal of creating an Italian-native resource that accurately represents Italian culture, we implemented a preliminary screening process to ensure that participants would provide answers from the point of view of an Italian person. For this data collection step, we accepted only annotators aged $25$ to $80$ who were born in Italy, spoke Italian as their first language, and had spent the majority of their first 18 years of life in Italy.
As with the question collection step, we used Google Forms (Figure 8) to provide the task to the annotators that was paid £$7$ per hour. Each form included $10$ videos, and for each video, the annotators were asked to answer $12$ questions [^12] Some annotators were given the $12$ Questions_A for that video and others the $12$-Questions_B group (one per category). The forms provided detailed guidelines on how to answer and specified the criteria for an acceptable response, which was necessary for the annotators to receive payment. Each response needed to be a complete answer (i.e., not short answers with missing noun / prepositional phrases), rely only on visible or deducible information from the video, and exclude any information derived from the video’s audio. Annotators were also encouraged to use their own world knowledge when interpreting the visual content.
Figure 8: Example of a Google Form used for collecting answers for each video
The initial acceptability of the task was determined by electing $2$ control questions from the $12$ semantic categories: the Out-of-scope and the Uncertainty question. Both questions had predefined acceptable answers, which made them easy to use for this purpose: the former required the negation of what asked in the question, while the latter required an expression of uncertainty in the response. To evaluate the acceptability of these control question answers, we employed the GPT-4o model in a few-shot prompting setting. During each annotation task, annotators worked with $10$ videos, each containing the $12$ questions,including the control questions. Annotation work was accepted only if the annotators correctly answered at least $90\%$ of the control questions (i.e., $18$ out of $20$ correct answers); otherwise, their submissions were rejected, and the task was reassigned to another annotator.
Our goal was to collect 8 different answers for each question to ensure not only accuracy but also variability in responses. To achieve this, we presented the same question to $8$ different annotators, collecting their individual answers. In total, $2,400$ questions were paired with 8 answers each, resulting in $19,200$ responses.
These responses then were features by a semi-automated review and validation process divided into two main steps, both facilitated by GPT-$4$o with few-shot prompting.
  1. 1.
    Semantic Consistency check. Each response was evaluated for semantic consistency with the corresponding question (Figure 9A). In cases where inconsistencies were detected, the responses were manually reviewed to determine whether the question should be re-answered by another annotator or the response could still be accepted. Real inconsistencies were found to be minimal (i.e., fewer than $100$ out of $19,200$ responses).
  2. 2.
    Contradiction test. By using GPT-4o OpenAI et al. (2024), we checked whether, within each pool of $8$ responses to the same question, any of the response contradicted the others (Figure 9B). We found that 90.25% of the 8-answer pools exhibit full agreement, as they do not contain any contradictions. The remaining 9.75% (234 cases) was manually reviewed by an additional annotator to resolve inconsistencies. The proportion of non-contradiction within the human annotated answers can also be considered as a rough proxy for Inter-annotator Agreement (IAA), since traditional IAA metrics are not easily applicable due to the open-ended nature of MAIA questions.
Figure 9: Prompts used for Answers’ semi-automatic validation. Prompts were provided with some examples (i.e., few-shot prompting) but here in the Figure only the main instructions are presented.
After completing these reviews, we obtained a dataset of $2,400$ questions, each paired with a pool of $8$ high-quality responses, resulting in a total of $19,200$ validated responses.
A.4 Answers Post-Processing
Since the data obtained up to this point will form the solid basis for the subsequent benchmarking phases (A.5 and A.7), we decided to set up a post-processing phase to assess and check that a certain degree of variability was ensured within each of the $2,400$ $8$-Answers pool. This work was carried out in several incremental stages, each of which involved an initial analysis of the potential redundancy of responses within the pool. Each of these analyses was then followed by a specific phase of semi-automatic expansion of these responses through the introduction of lexical variability.
  1. 1.
    Equal strings: a classic check to see if there were any pairs of equal answers. Table 6 shows the number of identified equal pairs divided by semantic category, highlighting a higher score for Out-of-scope Answers. Based on these results, GPT-4o was used in a Sentence Rephrasing task in a zero-shot prompting setting (Figure 10A). The only exception was the Out-of-scope category where all 1600 responses were manually edited.
  2. 2.
    Lexical Overlap: we tokenized each answer, checked the number of Type-Tokens shared between answer pairs within the pool of 8 and selected those that shared more than $50$% of them. This phase was followed for these selected items by an automatic Sentence Rephrasing task in which GPT-4o generated a new Answer with a Lexical overlap < $50$% with reference to the other $7$ Answers in the pool (Figure 10B). Table 7 shows the lexical overlap statistics before and after reformulation. The final average lexical overlap score within a pool($18.74$) represents the most balanced situation achievable. Indeed, it was noted that further attempts at lexical modifications within the pool would increase this percentage rather than make it lower. Finally, by focusing only on tokens represented by content words (i.e. adjectives, nouns, verbs and adverbs), it was noted that the lexical overlap index drops even further to $17.60$%.
Figure 10: Prompts used for Answers’ post-processing with GPT4-o
These two checks were finally followed by an analysis performed using the Type-Token Ratio (TTR) as a measure to quantify lexical richness (Table 7). The final results showed that, considering only content words, only $633$/$2400$ pools had a TTR < $0.5$ and that the average TTR per pool stands at $0.55$.
Category Equal Pairs
Causal 21
Counterfactual 2
Partial Implicit 133
Total Implicit 93
Uncertainty 112
Out-of-scope 414
Planning 2
Sentiment 9
Partial Spatial 77
Total Spatial 101
Temporal Duration 91
Partial Temporal 9
Tot. 1064
(2128 / 19200 Answers)
Table 6: Number of equal answers pairs in our
dataset before post-processing divided by semantic categories.
Before Rephrasing After Rephrasing
CW-only
Pool Avg. Lexical Overlap $22.95$% $18.74$% $17.60$%
Pool with Lexical Overlap > $50$% $11$ 00 00
Avg TTR for pool $0.50$ $0.55$
Pool with TTR < $50$% (Avg.) $1197$/$2400$ $633$/$2400$
Table 7: Lexical Overlap and TTR Statistics (Type-token) before and after Automatic Sentence Rephrasing with GPT-4o for $19,200$ Answers. Content-words (CW only) statistics are shown only for the After Rephrasing part.
After these steps, Figure 11 represents a valuable example of a high-quality question with a pool of $8$ answers in which we can find a significant level of lexical variability.
Figure 11: Example of a video with a Questions (Implicit category) and its 8-Answers pool.
A.5 True Statement Generation
After the collection and validation of the questions and answers, the next step involved the automatic generation of true statements (TS) which consists of descriptive declarative sentences aligned with the visual content of the videos. Just as a caption describes an image, a TS describes a video (or part of it) from different perspectives according to the specific semantic category.
For example, suppose a video shows a boy who is initially in a kitchen and while he is cooking he hears a loud noise and runs away to another room. An example of TS for the Spatial category could be:
Italian: Nel video il ragazzo è in cucina prima di scappare via.
English: In the video the boy is in the kitchen before running away.
In contrast a Causal TS:
In the scene the boy runs away because of the loud noise.
To create such statements, we used, as in the previous steps, GPT-4o (see Figure 12A for the prompt used). Using the information conveyed by the single question and its answer, we were able to combine them, generating $19,200$ true statements (TSs). As with the answers, the TS are also organised into $2,400$ pools of $8$ items. Each of these (TSs) in a pool will express the same event with different words and/or points of view.
A.6 True Statement post-processing
Just like with the pools of 8 responses, we also performed checks and semi-automatic interventions to ensure lexical variability within the 2400 pools of true statements (TS).
The steps followed were the same, except that for the TS we used the Type-Token Ratio (TTR) as an additional tool to enhance lexical variability within the pools. First, we identified $326$ pairs of identical TS and addressed this by regenerating the sentences using GPT-4o, as already done in LABEL:subsec:A-postprocessing, with the prompt shown in Figure 10A. Then, we analyzed the lexical overlap within the pools, focusing on statements with an overlap greater than $50$% ($354$ cases). This reduced the average overlap percentage of a pool from $39.34$% to $30.51$% (Type-Token) as shown in Table 8.
As a final step, we used the TTR to increase lexical richness within individual pools. We identified pools with a TTR < $0.5$ and replaced content words with synonyms to ensure that each TS contained different words. This approach allowed us to achieve a TTR of $0.50$ ($0.55$ when considering only content words), which is an excellent result given that, unlike responses, TS have a more fixed and repetitive structure (e.g., "In the video X happens" "The video shows X" etc.).
Before Rephrasing After Rephrasing
Type-Token CW-only Type-Token CW-only
Pool Avg. Lexical Overlap $39.34$% $38.04$% $30.51$% $26.81$%
Pool with Lexical Overlap > $50$% $354$ $344$ 00 00
Avg TTR for pool $0.37$ $0.41$ $0.50$ $0.55$
Pool with TTR < $50$% (Avg.) $2288$ $1964$ $1196$ $634$
Table 8: Lexical Overlap and TTR Statistics (Type-token and Content word-only) before and after Automatic Sentence Rephrasing with GPT-4o for 19200 TSs.
A.7 False Statement Generation
The automatic generation of true statements represents the final phase of data creation, ensuring all elements are ready to begin the actual benchmark task. Starting from the notion of the dichotomy of Caption and Foil —where a caption is a brief and correct descriptive sentence about a scene, and a foil is a minimally altered version of the caption that turns it into an incorrect description— we developed our False Statements (FSs).
The final goal of this work is to create minimal pairs for each semantic category, enabling controlled experiments and precise analysis of a model’s behavior with respect to the specific category.
As for TSs, the FSs were also automatically generated using $GPT-4o$. Figure 12B illustrates the prompt used during this phase, where the task was presented to the model using the Caption/Foil terminology rather than True/False Statements to enhance the effectiveness of the output. Given a semantic category and a caption, the task was to generate the corresponding foil by editing only the elements of the sentence related to that semantic category (e.g., If the category is Spatial, the prepositional phrase conveying spatial or location information is the block to be edited for creating the corrisponding foil).
Figure 12: Prompts used for True (A) and False (B) Statements generation with GPT-4o. Prompt B is representative of the 12 different prompts used to generate False Statements, each tailored to a specific semantic category.
A.8 False Statement Validation
As the last step of this work phase concerning the generation of data, we focused on the validation of the false statements we had just obtained. This review was characterised by the presence of two different semi-automatic checks based on the actual definition of foil: A description of the visual content minimally different from the caption, such that this difference falsifies its veracity. In both cases, GPT-4o was used as a model for the implementation of the task in a semi-automatic fashion.
  • Structural check: a minimal prompt (see Figure 13A) was constructed to check the correctness of a foil with respect to the related caption according to the specific category (e.g., if the category is Spatial, investigate whether, given a caption and a foil, the foil is correct with respect to the spatial information). Overall, $2.5$% of the foils generated turned out to be incorrect and then manually fixed. Table 9 shows for each semantic category the number of foils that resulted as incorrect according to GPT-4o, as opposed to those that actually were not and that were then manually corrected.
  • Contradiction test: based on the notion that the correct foil must be in contradiction with the relevant caption, we then ran an NLI task to identify cases of Entailment, Neutrality and Contradiction between our pairs of True and False Statements(Figure 13B). After a qualitative analysis of these cases, we realised that the cases identified by the model as Neutral (1287), were in fact correct and already partly defined a contradiction as in the next example.
    Italian:
    TS: I bambini toccano la lana per capirne la consistenza.
    FS: I bambini toccano la lana per capirne il colore.
    English:
    TS: The children touch the wool to check its texture.
    FS: The children touch the wool to check its colour.
    For this reason, the focus shifted only to the cases defined as Entailment ($170$). Within this group, only $93$ turned out to be true Entailments and they were subsequently manually modified.
This last revision allow us to obtain a packet of $19,200$ high-quality FFs with respect to its relative TSs.
Semantic Category FSs Incorrect FSs
GPT-4o Human
Causal 1600 80 54
Counterfactual 1600 61 15
Partial Implicit 1600 94 38
Total Implicit 1600 99 63
Uncertainty 1600 27 20
Out-of-scope 1600 48 9
Planning 1600 159 117
Sentiment 1600 10 7
Partial Spatial 1600 149 58
Total Spatial 1600 33 14
Temporal Duration 1600 33 29
Partial Temporal 1600 71 55
Tot. 19200 864 479
Table 9: Results of Semi-automatic False Statements Validation (Structural check)
Figure 13: Prompts used for False Statement Validation: Structural check (A) and Contradition Test (B)
Appendix B Experiments
This appendix section will contain additional details on our experimental settings, including a description of the Vision-Language Models used, as well as graphs and tables summarizing the results for Tasks 1 and 2 of MAIA. These supplementary materials provide a more comprehensive view of our findings, which could not be included in the main paper due to space limitations.
Note that in contrast to the initial experiments for creating and validating the synthetic data of the MAIA dataset, where we used OpenAI’s GPT-4o API, the experiments on MAIA were conducted using A100100100100 GPUs ($40$GB). Overall, the total computational budget was on the order of \$̃1,000\$ GPU hours.
B.1 Models tested
Vision-Language Models
We benchmarked four recent VLMs, sourced from the Hugging Face hub, representing state-of-the-art approaches in Vision-Language tasks. Both experiments and related quantitative evaluation have been done using lmms-eval (Zhang et al., 2024a), a framework for the evaluation of multimodal models.
InternVL2 Chen et al. (2024): $8$B parameter transformer-based multimodal model employing advanced cross-attention; pre-trained on large-scale image-text and video datasets for diverse multimodal tasks and instruction-tuned. It uses InternLM2.5 as open-sourced-$7$B parameter chat model.
Hugging Face model: OpenGVLab/InternVL2-8B
LLaVA-NeXT-Video Zhang et al. (2024b): $7$B parameter model built on the LLaVA framework, optimized for video understanding with mechanisms to capture temporal dynamics; fine-tuned on video instruction data. Base LLM: Vicuna-$7$B (v1.51.51.51.5).
Hugging Face model: llava-hf/LLaVA-NeXT-Video-7B-hf
LLaVa-OneVision Li et al. (2024a): $7$B parameter model that builds on the LLaVA framework with a Qwen2 LLM backbone to serve as a general-purpose vision-language assistant; pre-trained on extensive multimodal data to deliver robust cross-modal reasoning.
Hugging Face model: lmms-lab/llava-onevision-qwen2-7b-ov
Qwen2.5-VL Qwen et al. (2025): the latest $7$B parameter VLM of the Qwen family using the Qwen2.52.52.52.5 LLM decoder; key enhancements are related to grounding, working with longer videos and capturing events. It was pre-trained on comprehensive visual and textual datasets and fine-tuned for detailed, context-aware responses.
Hugging Face model: Qwen/Qwen2.5-VL-7B-Instruct
Unimodal models.
As described in Section 4, we used MAIA for benchmarking also some LLMs in a text-only evaluation. We used five open-weight LLMs which have shown good performance on a variety of tasks on Italian Magnini et al. (2025). For conducting these experiments, we used Minicons library Misra (2022), a high-level wrapper around Hugging Face for investigating predictive behavior of transformer models. Specifically, probabilities were computed adding a normalization parameter to take into account the different length of sentences in terms of tokens. Models used in the Hugging face Hub are: Llama-$3.1$ ($8$B-Instruct), LLaMAntino-$2$ ($7$B), LLaMAntino-$3$-ANITA ($8$B-Instruct), Gemma ($7$B) and Qwen2.52.52.52.5 ($7$B-Instruct).
B.2 Task1: Visual Statement Verification
As described in Section 4.4, the performance of the models in this task was compared against our most probable baseline. Figure 14 illustrates this relationship for each semantic category. Interestingly, text-only LLMs sometimes achieve higher accuracy in selecting the correct statement than VLMs, despite lacking visual input. For instance, in the counterfactual category the baseline values closely match those of the VLMs, while in the uncertainty category the baseline even outperforms both InternVL and LLaVA-NeXT-Video. Overall, the accuracy gap between our baseline and the VLMs is generally smaller than one might expect or theoretically anticipate. Table 10 presents the results of Task $1$, providing a detailed analysis for the Causal category, distinguishing between Explicit and Implicit. The same analysis is conducted in Table 13 for Task $2$ taking into account LLM’s judgements. Figure 15 represetns models’ fingerprint through MAIA’s reasoning categories, specifically illustrating models’ performance in the NLU Task 1 while Figure 16 reports performance on the same task with MAIA’s 8-answer pools.
Models Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Explicit Implicit Partial Total Partial Total Duration Partial
Unimodal 0.56 0.43 0.48 0.73 0.55 0.68 0.53 0.79 0.50 0.51 0.48 0.51 0.53 0.48
Black video InternVL2 0.68 0.61 0.67 0.88 0.89 0.69 0.61 0.96 0.52 0.59 0.54 0.55 0.67 0.60
Llava-Next-Video 0.50 0.48 0.50 0.52 0.48 0.51 0.52 0.45 0.50 0.51 0.51 0.49 0.51 0.48
Llava-oneVision 0.59 0.52 0.57 0.48 0.92 0.61 0.38 0.97 0.52 0.53 0.50 0.52 0.60 0.51
Qwen-2.5-VL 0.73 0.66 0.70 0.86 0.87 0.75 0.72 0.98 0.56 0.62 0.61 0.65 0.73 0.70
1-Frame InternVL2 0.75 0.77 0.82 0.87 0.69 0.71 0.77 0.89 0.65 0.80 0.65 0.82 0.67 0.69
Llava-Next-Video 0.61 0.67 0.70 0.76 0.49 0.63 0.75 0.40 0.59 0.65 0.59 0.70 0.56 0.53
Llava-oneVision 0.76 0.75 0.83 0.78 0.89 0.74 0.76 0.91 0.68 0.81 0.64 0.81 0.70 0.67
Qwen-2.5-VL 0.79 0.79 0.83 0.81 0.91 0.75 0.82 0.70 0.82 0.96 0.67 0.82 0.69 0.73
32-Frames InternVL2 0.79 0.82 0.84 0.85 0.77 0.75 0.84 0.84 0.75 0.83 0.69 0.86 0.66 0.76
Llava-Next-Video 0.52 0.54 0.57 0.57 0.42 0.57 0.61 0.32 0.52 0.59 0.52 0.56 0.51 0.48
Llava-oneVision 0.81 0.86 0.88 0.78 0.88 0.76 0.88 0.85 0.80 0.85 0.71 0.87 0.65 0.80
Qwen-2.5-VL 0.84 0.89 0.90 0.80 0.89 0.78 0.86 0.92 0.82 0.88 0.83 0.90 0.75 0.81
Table 10: Visual statement verification (Task $1$): accuracy of correct choices across reasoning categories with Causal sub-categories specification.
Figure 14: Accuracy of 4 VLMs compared with our most-probable baseline with reference to each category in MAIA
Figure 15: Task 1
Figure 16: Task 1 Aggregate
Models consistency.
In Task 1, model consistency in making uniform decisions within the same pool of True-False Statement pairs was investigated through a quantitative analysis, as outlined in Section 5.1. Moreover, Table 12 reports the consistency of models across the same pool of TF-FS pairs in Task $1$. As we can see, imposing a fair and rigid evaluation regime, viz. evaluating a model positively only when consistent across datapoints (total agreement), the 1-frame bias is mitigated for almost all models. The most consistent model is Qwen2.5-VL, while LLaVA-NeXT-Video performs best when a simple majority (at least $4$/$8$ consistent decisions) is considered.
Models Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Partial Total Partial Total Duration Partial
Unimodal 0.56 0.45 0.73 0.55 0.68 0.53 0.79 0.50 0.51 0.48 0.51 0.53 0.48
Black video InternVL2 0.68 0.64 0.88 0.89 0.69 0.61 0.96 0.52 0.59 0.54 0.55 0.67 0.60
Llava-Next-Video 0.50 0.49 0.52 0.48 0.51 0.52 0.45 0.50 0.51 0.51 0.49 0.51 0.48
Llava-oneVision 0.59 0.55 0.48 0.92 0.61 0.38 0.97 0.52 0.53 0.50 0.52 0.60 0.51
Qwen-2.5-VL 0.73 0.68 0.86 0.87 0.75 0.72 0.98 0.56 0.62 0.61 0.65 0.73 0.70
1-Frame InternVL2 0.75 0.80 0.87 0.69 0.71 0.77 0.89 0.65 0.80 0.65 0.82 0.67 0.69
Llava-Next-Video 0.61 0.68 0.76 0.49 0.63 0.75 0.40 0.59 0.65 0.59 0.70 0.56 0.53
Llava-oneVision 0.76 0.79 0.78 0.89 0.74 0.76 0.91 0.68 0.81 0.64 0.81 0.70 0.67
Qwen-2.5-VL 0.79 0.81 0.81 0.91 0.75 0.82 0.70 0.82 0.96 0.67 0.82 0.69 0.73
32-Frames InternVL2 0.79 0.83 0.85 0.77 0.75 0.84 0.84 0.75 0.83 0.69 0.86 0.66 0.76
Llava-Next-Video 0.52 0.56 0.57 0.42 0.57 0.61 0.32 0.52 0.59 0.52 0.56 0.51 0.48
Llava-oneVision 0.81 0.87 0.78 0.88 0.76 0.88 0.85 0.80 0.85 0.71 0.87 0.65 0.80
Qwen-2.5-VL 0.84 0.89 0.80 0.89 0.78 0.86 0.92 0.82 0.88 0.83 0.90 0.75 0.81
Table 11: Visual Statement Verification (Task $1$): accuracy of correct choices across reasoning categories.
Models Total Agreement Majority Agreeement
TS FS TS FS
Black video InternVL 17.83% 0.58% 63.79% 17.80%
LLaVa-neXT-Video 0.21% 0.29% 63.20% 36.30%
LLava-oneVision 13.91% 1.71% 57.20% 27.18%
Qwen2.5-VL 22.25% 0.54% 66.25% 10.96%
1-Frame InternVL 27.58% 0.70% 60.42% 11.30%
LLaVa-neXT-Video 8.30% 0.96% 68.70% 22.04%
LLava-oneVision 31.55% 0.71% 56.62% 11.12%
Qwen2.5-VL 35.92% 0.25% 55.33% 8.50%
32-Frame InternVL 31.25 % 0.30% 60.04% 8.41%
LLaVa-neXT-Video 2.79% 1.46% 62.25% 33.5%
LLava-oneVision 37.71% 0.21% 54.33% 7.75%
Qwen2.5-VL 43.62% 0.21% 50.92% 5.25%
Table 12: Task 1: consistency of TS and FS across the 8-answer pool. Total agreement is 8/8, majority is between 4 and 7. There are 00 cases with agreement < 4.
We further examined the accuracy variability across the 8 pools, with Figure 17 providing a detailed examination of trends within each category. This analysis revealed notable fluctuations in accuracy for certain categories such as Uncertainty, Spatial, and Temporal categories, pointing to a lack of stability and highlighting the complexity of tasks . For instance, some models, like LLaVA-NeXT-Video, exhibited more significant variability, showcasing their challenge in handling these categories. Despite these fluctuations, Qwen2.5-VL emerged as the top performer overall, maintaining the most consistent accuracy across pools. Similar conclusions can be reached through visual inspection of Figures 15 and 16.
Figure 17: Performance trends of the four models. Each plot shows the accuracy trend as a function of the 8 different pools.
B.3 Task2
We employed two evaluation approaches for the generation task: similarity-based metrics and the LLM-as-a-judge approach. The latter proved to be the most suitable for our purposes, given its flexibility and ability to handle open-ended responses effectively. The use of 8 reference answers was instrumental in this process. Inspired by the excellent work of Lee et al. (2024a) and Mañas et al. (2024), we leveraged these references to assess model-generated responses without enforcing strict syntactic similarity. This approach allows for a more nuanced evaluation, prioritizing semantic alignment over surface-level structure. Notably, these reference answers are slightly different but all correct, capturing a range of valid perspectives. This ensures a more robust and reliable evaluation, as it accounts for the natural variability in human responses and the multiple ways a model might correctly express the same information. To ensure further robustness, we adopted GPT-4o as our evaluation model, prompting it to determine correctness by checking whether a candidate answer aligns with at least one reference. Most importantly, we only proposed this method after verifying its high correlation with two human annotators (Fleiss’ Kappa: 0.82), ensuring its reliability. For transparency, Figure 18 represent the prompt used for the automatic evaluation of VLMs’ answers. Figure 19 represetns models’ fingerprint through MAIA’s reasoning categories, specifically illustrating models’ performance in the NLG Task 2.
Figure 18: Prompt used for automatic evaluation of VLMs’ answers in Task2 (i.e., for LLM-as-a-judge evaluation metric)
Models Avg. Causal Counterfactual Out-of-Scope Planning Sentiment Uncertainty Implicit Spatial Temporal
Explicit Implicit Partial Total Partial Total Duration Partial
Black video InternVL2 0.37 0.41 0.43 0.60 0.30 0.68 0.43 0.08 0.21 0.23 0.36 0.25 0.55 0.25
Llava-Next-Video 0.27 0.40 0.47 0.47 0.30 0.40 0.33 0.51 0.21 0.12 0.16 0.04 0.12 0.16
Llava-oneVision 0.40 0.60 0.73 0.68 0.29 0.60 0.60 0.28 0.26 0.24 0.36 0.18 0.37 0.23
Qwen-2.5-VL 0.35 0.35 0.37 0.69 0.08 0.54 0.74 0.23 0.24 0.19 0.30 0.19 0.41 0.20
1-Frame InternVL2 0.44 0.48 0.67 0.65 0.21 0.65 0.60 0.10 0.33 0.38 0.39 0.47 0.53 0.35
Llava-Next-Video 0.32 0.24 0.36 0.56 0.20 0.46 0.60 0.29 0.18 0.32 0.24 0.27 0.20 0.19
Llava-oneVision 0.50 0.56 0.62 0.78 0.15 0.66 0.74 0.37 0.39 0.54 0.39 0.51 0.54 0.33
Qwen-2.5-VL 0.51 0.50 0.62 0.76 0.32 0.60 0.79 0.40 0.35 0.50 0.38 0.53 0.55 0.38
32-Frames InternVL2 0.49 0.50 0.58 0.68 0.28 0.64 0.62 0.11 0.45 0.48 0.47 0.51 0.57 0.46
Llava-Next-Video 0.33 0.33 0.42 0.38 0.16 0.42 0.48 0.27 0.24 0.37 0.29 0.39 0.32 0.29
Llava-oneVision 0.53 0.63 0.71 0.79 0.11 0.65 0.79 0.23 0.55 0.56 0.51 0.62 0.40 0.46
Qwen-2.5-VL 0.61 0.68 0.75 0.80 0.43 0.60 0.85 0.55 0.55 0.60 0.50 0.70 0.54 0.53
Table 13: Open-ended Visual Question Answering (Task $2$): accuracy of correct answers with LLM-as-a-judge with Causal sub-categories specification.
Figure 19: Task 2
B.4 Case study: Multimodal Hallucinations
As part of a preliminary error analysis focusing on the Out-of-Scope category, we observed that a significant portion of the errors made by the model could be attributed to multimodal hallucinations. Particularly interesting was the discovery of counterintuitive clashes between the two tasks in our benchmark. In several instances, the model successfully solved Task 1 (i.e., True Statement Selection), in some cases achieving also full consistency within the pools (e.g., 8/8 correct selections), yet failed Task 2 (i.e., open-ended NLG), generating hallucinated content. For example, given a video in which a young girl is frightened by an insect in her home and seeks her mother’s help to remove it, our best-performing model (i.e., Qwen2.5VL) consistently selected the correct true statement in all eight pairs [^13] True Statements are the first in each pair:
  1. 1.
    • In the scene there is no dog at the door
    • In the scene there is a dog at the door
  2. 2.
    • In the movie there are no animals at the door
    • In the movie there are animals at the door
  3. 3.
    • In the movie there is no doggie at the doorway
    • In the movie there is a doggie at the doorway
  4. 4.
    • No pets are seen in front of the door
    • Some Pets are seen in front of the door
  5. 5.
    • There are no dogs at the door in the video
    • There are dogs at the door in the video
  6. 6.
    • No dog appears in the entrance area
    • A dog appears in the entrance area
  7. 7.
    • No pets are visible in the video near the front door of the house
    • A pet is visible in the video near the front door of the house
  8. 8.
    • In the video clip there is no dog at the door
    • In the video clip there is a dog at the door
However, when prompted in Task 2 with the question What color is the dog at the door?, the model hallucinates by answering The dog at the door is black, despite the fact that no dog is present in the video. This case highlights a curious but also dangerous misalignment between the model’s apparent ability to correctly perform discriminative reasoning in a multiple-choice setting and its failure to accurately generate grounded content. It also emphasizes a lack of robustness in the model’s competencies. These findings underscore the need for a more in-depth error analysis, as the overall results (see Tables 2 and 3) suggest that, for specific reasoning categories, performance in Task 1 may not reliably predict success in Task 2, with notable performance unbalances often occurring (e.g., uncertainty reasoning category).